# Choosing observables that capture critical slowing down before tipping points: A Fokker-Planck operator approach

### v1

@@ III.3 Eigenfunction reconstruction from diffusion maps and observables for early-warning signals
Third, a physical observable can be constructed by expressing $\xi_{1}$ as function of (possibly a subset of) the state variables. $\xi_{1}$ is estimated for an observational time slice closest to the TP and the values of $\xi_{1}$ at the data points (i.e. the entries of the eigenvectors of the Markov matrix $M$) are fit to a suitable function, e.g., a polynomial. The fitted function can then be evaluated as an observable for any data sets further away from the TP. In Fig. 6e-g we show polynomial fits to $\xi_{1}$ estimated from simulation data of the DW2 model. The directionality of the level sets of the observable functions is consistent with the direction of the edge state, as expected. Far from the bifurcation (assuming low noise) the dynamics samples only the relatively flat part of $\xi_{1}$ far from the saddle point, provided the noise is sufficiently small, as shown in Fig. 6e. Still, the fitted function shows a more rapid decrease towards the saddle point, thus indicating that it already carries the crucial information for detecting CSD. When close to the TP, non-linear functions tend to be required for an adequate fit of the $\xi_{1}$ data (Fig. 6f). However, linear fits preserve the directionality of the edge state well (Fig. 6g). Using the fits estimated from data sampled close to the TP (Fig. 6f,g), we again see that the variability of the values of $\xi_{1}$ decreases when evaluated for data sample further away from the TP (Fig. 6c,d). The non-linear observable shows a significantly stronger change in variance compared to the computed linear one (Fig. 6b).
@@
Similar results are obtained when applying the method to a slightly more complex non-gradient system, which is summarized in the following for a four-dimensional conceptual model of the AMOC [41]. The variables ($\{S_{N},S_{T},S_{S},S_{I}\}$) are the average salinities in four boxes of the global ocean (see Sec. A for the equations). The large difference in the box volumes gives a time scale separation, with $S_{S}$ and $S_{I}$ being the slowest variables, and $S_{N}$ being the fastest. The model is bistable for a range of the control parameter $H$ from $H\approx 0.04$ until the bifurcation at $H\approx 0.3214$, where the stable fixed point corresponding to a present-day AMOC disappears.
Using simulation data restricted to the present-day AMOC state, far from the bifurcation $\xi_{1}$ represents a correlated relaxation mode in the slow variables $S_{S}$ and $S_{I}$. $\xi_{2}$ is a quadratic function of $\xi_{1}$, and $\xi_{3}$ is independent and strongly correlated with $S_{N}$ only (Fig. 7a,b). When approaching the TP, the slow mode corresponding to relaxation along the direction of the edge state emerges due to CSD, and eventually rises to the position of $\xi_{1}$, with $\xi_{2}$ ($\xi_{3}$) being a quadratic (cubic) function thereof (Fig. 7c,d). $\xi_{1}$ exhibits a strong non-linear anti-correlation with $S_{N}$ and a non-linear positive correlation with $S_{T}$ (see also [15]). A good physical observable representing $\xi_{1}$ is found by considering polynomial functions in a projected space of a subset of (e.g. two) variables. The best cubic polynomial fit of $\xi_{1}$ (estimated at $H=0.315$) is a function of $S_{N}$ and $S_{T}$. The resulting function shows a monotonic, non-linear increase from the base state towards the edge state, while being essentially flat in other directions (Fig. 7e). This renders the observable very sensitive to excursions towards the edge state, and thus ideal for EWS.
@@ III.4 Observables and extrapolation of tipping times
So far we discussed the construction of observables that show a large increase in variance (and also autocorrelation) as a result of CSD, which is a qualitative indication that the system moves towards a TP. One may go further and attempt a quantitative prediction of the expected time of tipping by extrapolating the CSD signal in observational time series, as was done in the context of real-world climate observations in [20, 24]. This works by assuming that an observed time series samples the critical dynamics by obeying the SNB normal form, implying that such a prediction is sensitive to the choice of observable.
Consider the general multi-dimensional system described by the coupled stochastic differential equations (1), where we now assume that the drift $a^{\gamma}(\mathbf{X},\mu)$ depends on a control parameter $\mu$ and the noise $\sigma_{\nu}^{\gamma}(\mathbf{X})=\sigma_{\nu}^{\gamma}$ is additive. If the system undergoes SNB the noise-driven dynamics is expected to become restricted to the vicinity of a one-dimensional center manifold, and is described by the normal form for a SNB
@@
where $\mu=0$ demarcates the bifurcation. Close to the fixed point, the system can be linearized and approximated by the Ornstein-Uhlenbeck process $d\tilde{x}_{t}=-\lambda\tilde{x}dt+\sigma dW_{t}$. Crucially, the linear restoring rate $\lambda$ is related to the control parameter with $\lambda=2\sqrt{\mu}$. Data that are sampled at small time intervals $\Delta_{t}$ can be approximated by an AR(1) process
@@
where $\epsilon_{k}$ are Gaussian random variables with variance $\sigma^{2}(2\lambda)^{-1}(1-e^{-2\lambda\Delta t})$. For an AR(1) process of this form the autocorrelation at lag 1 is given by $\rho_{1}=e^{-\lambda\Delta t}$. This means that from the above relation of $\lambda$ and the control parameter $\mu$ the autocorrelation tends to 1 at the SNB, and we can reconstruct $\mu$ from data by
@@
Thus, in a sliding window one can estimate $\rho_{1}$ as a function of time, and, assuming a linear trend in $\mu$, estimate with a linear fit to the function on the righthand-side at what time the control parameter will cross zero.
@@
Figure 8: Simulations of the DW1 model (16) with $\sigma=0.05$ and a linear increase of the control parameter within 500 time units from $\beta=-0.7$ past the bifurcation point $\beta_{c}=\sqrt{8/27}$ to $\beta=0.7$. a Time series of one simulation overlaid on the bifurcation diagram. b Reconstructed control parameter (Eq. 22) obtained by the lag-1 autocorrelation $\rho_{1}$, estimated at each time step (sample spacing of $\Delta t=0.05$ time units) for an ensemble of 15,000 simulations (black trajectory) until a cutoff time where 98% of the ensemble members have not tipped yet (evaluated by crossing a threshold of $x=0.2$). At the noise level $\sigma=0.05$, this is only very shortly before the bifurcation is reached. The 2% of realizations that tipped are removed. The solid lines are linear fits using different segments of the data, with the red (blue) line using the first half (last sixth) of the data. The crosses as well as dotted and dashed lines are for $\rho_{1}$ estimated in a moving window of length 40.
@@
If only given data from an arbitrary scalar observable of the system, this extrapolation to the time of tipping can fail since the observable need not obey the SNB normal form, or may only approximately do so when arbitrarily close to the bifurcation. In fact, even for bi-stable systems with only one variable (where the question of observable is obsolete), the scaling $\mu\propto[\ln(\rho_{1}(X))]^{2}$ according to the saddle-node normal form only applies when close to the bifurcation. For the DW1 model (16) under a linear change in time of the control parameter $\mu$ the function $[\ln(\rho_{1}(X))]^{2}$ is a convex function of time (Fig. 8). Thus, the TP, which occurs at $t\approx 420$, would be predicted too early at $t\approx 390$ (red line in Fig. 8b) when the linear extrapolation is performed based on data not close enough to the bifurcation. An estimation of the autocorrelation in a moving window introduces a further bias towards a later estimated tipping, because the autocorrelation is underestimated by removing some correlation during the necessary step of detrending within each window.
@@
Moreover, apart from these biases, as more dimensions are involved the prediction depends on the choice of observable. For the DW2 model (19), the variable $x$ can give an accurate prediction when data are available close enough to the TP (Fig. 9b). In contrast, $y$ initially shows a quasi-linear relation of $\mu$ and $[\ln(\rho_{1}(X))]^{2}$, but then a much steeper relationship closer to the bifurcation (Fig. 9c). Extrapolating based on the initial slope would lead to an estimated tipping time that is far too late. Especially unsuited observables exist, such as $O(x,y)=x-y$, where no tipping can be predicted before a noise-induced transitions would occur (Fig. 9d). In contrast, $O(x,y)=x+y$ is very closely aligned with the direction of the edge state [15] and permits to predict the time of tipping accurately for data sufficiently close to the bifurcation (Fig. 9e). Finally, the non-linear observable obtained from the DM approximation to the first subdominant backward eigenfunction (Fig. 6f) is most accurate, even when evaluated at data far from the bifurcation (Fig. 9f).
@@
Figure 9: Same as Fig. 8, but for simulations with the DW2 model (19) with $\sigma_{x}=\sigma_{y}=0.05$, where the parameter $e$ is increased linearly past the bifurcation. Shown in a is a single example trajectory overlaid on the bifurcation diagram. Panels b-f show the evolution of the quantity $[\ln(\rho_{1}(O))]^{2}(2\Delta t)^{-2}$ for different observables $O(x,y)$, along with linear fits to different parts of the time series.
@@ III.5 Application to tipping points in a global ocean model
As the final result, we show our method is capable of successfully detecting TPs in a high-dimensional system. We consider the global ocean model Veros [42], which shows a TP of the AMOC from its present-day state to a collapsed state as a result of increasing meltwater input to the North Atlantic. Veros is a primitive-equation finite-difference ocean model forced with a fixed atmospheric climatology, and discretized on a grid of 40 latitudinal and 90 longitudinal grid points, as well as 40 depth levels. This is a coarse-resolution setup, but it enables long steady-state simulations to get good statistics beneficial for our feasibility study. As a dynamical system, the model possesses almost one million degrees of freedom. For more details on the model, see [42, 43, 44]. The meltwater input $F$ is the control parameter, and the stability landscape with respect to $F$ (computed in [44]) is shown in Fig. 10. There are several branches of attractors with an AMOC similar to present-day, but these collapse at a high freshwater forcing of $F\approx 0.36$. After this TP, there remain only attractor branches with a collapsed AMOC.
@@
We use four 33,000-year long simulations (after removal of a transient for equilibration) performed at four, fixed values of $F$ leading up to the TP, and sampled as 5-year averages of the state variables. These are referred to as simulations I to IV, see Fig. 10. Surface temperature and salinity noise forcing drives fluctuations of the system around its deterministic attractors (for more details see [15]), which otherwise feature relatively small-amplitude chaotic oscillations [44]. The system without noise forcing has been investigated previously to determine an edge state on the separatrix of the present-day and collapsed AMOC regimes [45]. By analyzing its mean climatology it was found that the edge state distinguishes itself most strongly from the mean states on the attractors in terms of its fresh and cold deep Atlantic. This is the “fingerprint” of the edge state. Subsequently, it was argued that increased fluctuations towards the edge state as a result of CSD should be most prominent in variables quantifying this fingerprint [15]. Indeed, only a very small subset of all d.o.f, coinciding exactly with the variables describing the deep ocean fingerprint, shows significant variance increase prior to the AMOC collapse [15]. The variable that was found to exhibit the largest increase in variance is shown in Fig. 11a-d across the four data sets. In the following, we show that similar (if not better) results can be obtained with our the DM method which does not require prior knowledge of the edge state or a brute-force search across all d.o.f (risking false positives). Instead, only observational data close to the TP is required.
@@
It would be feasible to compute the DM distance Kernel in the full space of the three-dimensional fields, perhaps after a weighting of the different physical units (temperature, salinity, density and velocity). But for simplicity we perform an initial dimensionality reduction, by averaging the salinity, temperature and density fields over boxes covering the entire ocean at different depths, and by summarizing the strength of the main ocean currents in terms of the spatial maxima of the meridional and barotropic stream functions (see [45]). This yields time series of 83 variables covering most important aspects of the model state. Fig. 11a-h shows time series of two of these variables at four different values of $F$, with the mean removed. Before applying the DM algorithm, we normalize all variables to have unit variance. A bandwidth of $\epsilon=35$ for the similarity Kernel (11) was found to be optimal to resolve the data manifold at all parameter values $F$ without being influenced by outliers, the largest of which are removed beforehand (see Sec. II).
Table 1: Spearman correlation of the top five Veros features with the first two diffusion coordinates $\xi_{1}$ and $\xi_{2}$, where the $\xi_{n}$ were estimated at the four different values of the control parameter $F$. In bold text are key variables (’rho’ refers to density) related to the mode of cold and fresh excursions in the deep north and tropical Atlantic directed towards the edge state. In italic are key variables related to the mode of fast cold excursions in the Southern ocean, which strongly increase the density in large parts of the deep ocean.
@@
A high-dimensional feature set gives much freedom in designing physical observables for EWS from $\xi_{1}$. The first option is again to use the Nyström extension based on all variables and interpolate the function $\xi_{1}$ to observations further back in time, in order to find evidence for increased fluctuations in the critical mode. Next, one may fit $\xi_{1}$ as a linear or non-linear function of the variables. In high dimensions it is sensible to only consider a subset of variables to find a parsimonious observable with the best signal-to-noise ratio when applied for EWS. We leave a treatment of this statistical optimization problem for future research, and consider here three simple examples. First, simply take the variable with highest correlation to $\xi_{1}$ as observable. This is NA deep ocean salinity (Fig. 11a-d), which indeed has the highest increase in variance of all individual features [15], increasing by a factor of 12.23 when going from $F=0.2957$ to $F=0.3557$. Second, fit $\xi_{1}$ to a linear combination of the two best features in Tab. 1 (at $F=0.3557$). This yields a variance increase by a factor of 14.21. Third, fit $\xi_{1}$ to a cubic polynomial of two variables, which are chosen as the two features among the best four (Tab. 1) that represent two different ocean sectors, thus giving some degree of spatial independence. This observable captures the directionality of the edge state well (Fig. 12), and shows an increase in variance by a factor of 12.60.
@@
Finally, $\xi_{1}$ can be mapped back into the physical space of full dimension, whereafter spatio-temporal anomalies from different time periods can be projected onto the mode. In particular, we propose to obtain the physical representation of this mode by averaging over the time points where $\xi_{1}$ is extremized. For instance, one can choose the data points with the top 5% largest and 5% smallest values of $\xi_{1}$. Averaging independently over these two sets of data points defines a positive and a negative phase of the mode. By taking the difference of the positive and negative patterns (or vice versa) we obtain a pattern that describes the mode as a whole and that we can project data onto. This may be viewed as linear approximation that interpolates the physical mode linearly as a function of the value of $\xi_{1}$. Fig. 13 gives a comparison of the first subdominant modes extracted in this way by extremizing $\xi_{1}$ far from (Fig. 13a,d) and close to the TP (Fig. 13b,e). The modes are projected down to the two-dimensional physical space of vertically averaged deep ocean of temperature and salinity. Far from the TP, the mode is characterized by a global cooling of the deep ocean initiated by abrupt (decadal-scale) cooling events in the Southern Ocean (seen as spikes of density increase in Fig. 11e-h), which are excited by noise in the multistability regime, as discussed in [15]. The resulting dense deep ocean water spreads throughout the deep ocean, before the signal decays on a multi-centennial time scale. There is no significant salinity signal.
@@
While these are encouraging results, in practice there is a risk of false positives and negatives, since here the critical physical mode only appears in its correct place at $\xi_{1}$ when already quite close to the TP. One needs to verify that the leading mode captured by $\xi_{1}$ is likely a critical mode. The guiding principle should be to look how strongly fluctuations in the leading diffusion coordinate, estimated at the current time slice, have increased compared to data slices back in time, and then set a level of statistical significance based on a reference period. In the Veros data one can in this way rule out the initially leading mode (Fig. 13a,d), because its variability does not keep increasing towards the TP (Fig. 11f-h), and the associated excursions do not last longer. But in principle there remains a chance for false positives of a new time scale separation would arise upon change of a control parameter for reasons not related to a TP.

### v2

@@ III.3 Eigenfunction reconstruction from diffusion maps and observables for early-warning signals
Third, an explicit expression of a physical observable can be constructed from $\xi_{1}$ as function of (possibly a subset of) the state variables. $\xi_{1}$ is estimated for an observational time slice closest to the TP and the values of $\xi_{1}$ at the data points (i.e. the entries of the eigenvectors of the Markov matrix $M$) are fit to a suitable function, e.g., a polynomial. The fitted function can then be evaluated for any data sets further away from the TP. In Fig. 6e-g we show polynomial fits to $\xi_{1}$ estimated from simulation data of the 2-d double-well. The directionality of the level sets is consistent with the direction of the edge state, as expected. Far from the bifurcation (assuming low noise) the dynamics samples only the relatively flat part of $\xi_{1}$ far from the saddle point, provided the noise is sufficiently small, as shown in Fig. 6e. Still, the fitted function shows a more rapid decrease towards the saddle point, thus indicating that it already carries the crucial information for detecting CSD. When close to the TP, non-linear functions tend to be required for an adequate fit of the $\xi_{1}$ data (Fig. 6f). However, linear fits preserve the directionality of the edge state well (Fig. 6g). Using the fits estimated from data sampled close to the TP (Fig. 6f,g), we again see that the variability of the values of $\xi_{1}$ decreases when evaluated for data sample further away from the TP (Fig. 6c,d). The non-linear observable shows a significantly stronger change in variance compared to the linear one (Fig. 6b).
@@
Similar results are obtained when applying the method to a slightly more complex non-gradient system, a four-dimensional conceptual model of the AMOC [49], which is summarized in the following. The variables ($\{S_{N},S_{T},S_{S},S_{I}\}$) are the average salinities in four boxes of the global ocean (see Sec. B for the equations). Large differences in the box volumes gives a time scale separation, with $S_{S}$ and $S_{I}$ being the slowest variables, and $S_{N}$ being the fastest. The model is bistable for a range of the control parameter $H$ from $H\approx 0.04$ until the bifurcation at $H\approx 0.3214$, where the stable fixed point corresponding to a present-day AMOC disappears.
We use simulation data restricted to the present-day AMOC state. Far from the bifurcation, $\xi_{1}$ represents a correlated relaxation mode in the slow variables $S_{S}$ and $S_{I}$. $\xi_{2}$ is a quadratic function of $\xi_{1}$, and $\xi_{3}$ is independent and strongly correlated with $S_{N}$ only (Fig. 7a,b). When approaching the TP, the slow mode corresponding to relaxation along the direction of the edge state emerges due to CSD, and eventually rises to the position of $\xi_{1}$, with $\xi_{2}$ ($\xi_{3}$) being a quadratic (cubic) function thereof (Fig. 7c,d). $\xi_{1}$ exhibits a strong non-linear anti-correlation with $S_{N}$ and a non-linear positive correlation with $S_{T}$ (see also [15]). A good physical observable representing $\xi_{1}$ is found by considering polynomial functions in a projected space of a subset of (e.g. two) variables. The best cubic polynomial fit of $\xi_{1}$ (estimated at $H=0.315$) is a function of $S_{N}$ and $S_{T}$. The resulting function shows a monotonic, non-linear increase from the base state towards the edge state, while being essentially flat in other directions (Fig. 7e). This renders the observable very sensitive to excursions towards the edge state, and thus ideal for EWS.
@@ III.4 Observables and extrapolation of tipping times
So far we discussed the construction of observables that show a large increase in variance (and also autocorrelation) as a result of CSD, which is a qualitative indication that the system moves towards a TP. One may go further and attempt a quantitative prediction of the expected time of tipping by extrapolating the CSD signal in observational time series, as was done in the context of real-world climate observations in [20, 24]. For this one needs to assume that an observed time series samples the critical dynamics by obeying the SNB normal form. This implies that such a prediction is sensitive to the choice of observable, as shown in the following.
Consider the general multi-dimensional system described by the coupled stochastic differential equations (1), where we now assume that the drift $b^{\gamma}(\mathbf{X},\mu)$ depends on a control parameter $\mu$ and the noise $\sigma_{\nu}^{\gamma}(\mathbf{X})=\sigma_{\nu}^{\gamma}$ is additive. If the system undergoes a SNB, the noise-driven dynamics is expected to become restricted to the vicinity of a one-dimensional center manifold, and is described by the normal form
@@
where $\mu=0$ demarcates the bifurcation. Close to the fixed point, the system can be linearized and approximated by the Ornstein-Uhlenbeck process $d\tilde{x}_{t}=-\lambda\tilde{x}dt+\sigma dW_{t}$. Crucially, the linear restoring rate $\lambda$ is related to the control parameter with $\lambda=2\sqrt{\mu}$. Data sampled at small time intervals $\Delta_{t}$ can be approximated by an AR(1) process
@@
where $\epsilon_{k}$ are Gaussian random variables with variance $\sigma^{2}(2\lambda)^{-1}(1-e^{-2\lambda\Delta t})$. For this process the autocorrelation at lag 1 is given by $\rho_{1}=e^{-\lambda\Delta t}$. Since $\lambda=2\sqrt{\mu}$, we can reconstruct $\mu$ from data by
@@
Thus, in a sliding window one can estimate $\rho_{1}$ as a function of time, and, assuming a linear trend in $\mu$, estimate with a linear fit to the function on the righthand-side at what time the control parameter will cross zero, which is exactly when the autocorrelation tends to 1 at the SNB.
@@
Figure 8: Simulations of the 1-d double well (16) with $\sigma=0.05$ and a linear increase of the control parameter within 500 time units from $\beta=-0.7$ past the bifurcation point $\beta_{c}=\sqrt{8/27}$ to $\beta=0.7$. a Time series of one simulation overlaid on the bifurcation diagram. b Reconstructed control parameter (22) obtained by the lag-1 autocorrelation $\rho_{1}$, estimated at each time step ($\Delta t=0.05$ time units) for an ensemble of 15,000 simulations (black trajectory) until a cutoff time where 98% of the ensemble members have not tipped yet (evaluated by crossing a threshold of $x=0.2$). At the noise level $\sigma=0.05$, this is only very shortly before the bifurcation is reached. The 2% of realizations that tipped are removed. The solid lines are linear fits using different segments of the data, with the red (blue) line using the first half (last sixth) of the data. The crosses as well as dotted and dashed lines are for $\rho_{1}$ estimated in a moving window of length 40.
@@
If given data from an arbitrary scalar observable of the system, this extrapolation to the time of tipping can fail since the observable need not obey the SNB normal form, or may only approximately do so when arbitrarily close to the bifurcation. In fact, even for univariate bi-stable systems (where the question of observable is obsolete), the scaling $\mu\propto[\ln(\rho_{1}(X))]^{2}$ according to the saddle-node normal form only applies when close to the bifurcation. For the 1-d double-well (16) under a linear change in time of the control parameter $\mu$ the function $\ln(\rho_{1}(X))]^{2}$ is convex with respect to time (Fig. 8). Thus, the TP, which occurs at $t\approx 420$, would be predicted too early at $t\approx 390$ (red line in Fig. 8b) when the linear extrapolation is performed based on data not close enough to the bifurcation. An estimation of the autocorrelation in a moving window introduces a further bias towards a later estimated tipping, because the autocorrelation is underestimated by removing some correlation during the necessary step of detrending within each window.
@@
As more dimensions are involved the prediction depends on the choice of observable. For the 2-d double-well (19), the variable $x$ can give an accurate prediction when data are available close enough to the TP (Fig. 9b). In contrast, $y$ initially shows a quasi-linear relation of $\mu$ and $\ln(\rho_{1}(X))]^{2}$, but then a much steeper relationship closer to the bifurcation (Fig. 9c). Extrapolating of the initial slope would lead to a tipping time estimate that is far too late. Especially unsuited observables exist, such as O(x,y)=xyO(x,y)=x-y, where no tipping can be predicted before a noise-induced transitions would occur (Fig. 9d). In contrast, $O(x,y)=x+y$ is very closely aligned with the direction of the edge state [15] and permits to predict the time of tipping accurately for data sufficiently close to the bifurcation (Fig. 9e). The non-linear observable obtained from the DM approximation to the first subdominant backward eigenfunction (Fig. 6f) is most accurate, even when evaluated at data far from the bifurcation (Fig. 9f).
@@
Figure 9: Same as Fig. 8, but for simulations with the 2-d double-well (19) with $\sigma_{x}=\sigma_{y}=0.05$, where the parameter $e$ is increased linearly past the bifurcation. Shown in a is a single example trajectory overlaid on the bifurcation diagram. Panels b-f show the evolution of the quantity $\ln(\rho_{1}(O))]^{2}(2\Delta t)^{-2}$ for different observables $O(x,y)$, along with linear fits to different parts of the time series.
@@ III.5 Application to tipping points in a global ocean model
As the final result, we show our method is capable of successfully detecting TPs in a high-dimensional system. We consider the global ocean model Veros [50], which shows a TP of the AMOC from its present-day state to a collapsed state as a result of increasing meltwater input to the North Atlantic. Veros is a primitive-equation finite-difference ocean model forced with a fixed atmospheric climatology, and discretized on a grid of 40 latitudinal and 90 longitudinal grid points, as well as 40 depth levels. This is a coarse-resolution setup, but it enables long steady-state simulations to get good statistics beneficial for our feasibility study. As a dynamical system, the model possesses almost one million degrees of freedom. For more details on the model, see [50, 51, 52]. The meltwater input $F$ is the control parameter, and the stability landscape with respect to $F$ (computed in [52]) is shown in Fig. 10. There are several branches of attractors with an AMOC similar to present-day, but these collapse at a high freshwater forcing of $F\approx 0.36$. After this TP, there remain only attractors with a collapsed AMOC.
@@
We use four 33,000-year long equilibrated simulations performed at four, fixed values of $F$ leading up to the TP, and sampled as 5-year averages of the state variables. These are referred to as simulations I to IV, see Fig. 10. Additive surface temperature and salinity noise drives fluctuations of the system around its deterministic attractors (for more details see [15]), which otherwise feature relatively small-amplitude chaotic oscillations [52]. The system without noise forcing has been investigated previously to determine an edge state on the separatrix of the present-day and collapsed AMOC regimes [53]. By analyzing its mean climatology it was found that the edge state distinguishes itself most strongly from the mean states on the attractors in terms of its fresh and cold deep Atlantic. This is the “fingerprint” of the edge state. Subsequently, it was argued that increased fluctuations towards the edge state as a result of CSD should be most prominent in variables quantifying this fingerprint [15]. Indeed, only a very small subset of all d.o.f, coinciding exactly with the variables describing the deep ocean fingerprint, shows significant variance increase prior to the AMOC collapse [15]. The variable that was found to exhibit the largest increase in variance is shown in Fig. 11a-d across the four data sets. In the following, we show that similar (if not better) results can be obtained with our the DM method which does not require prior knowledge of the edge state or a brute-force search across all d.o.f (risking false positives). Instead, only observational data close to the TP is required.
@@
It would be feasible to compute the DM distance Kernel in the full space of the three-dimensional fields, perhaps after a weighting of the different physical units (temperature, salinity, density and velocity). But for simplicity we perform an initial dimensionality reduction, by averaging the salinity, temperature and density fields over boxes covering the entire ocean at different depths, and by summarizing the strength of the main ocean currents in terms of the spatial maxima of the meridional and barotropic stream functions (see [53]). This yields time series of 83 variables covering most important aspects of the model state. Fig. 11a-h shows time series of two of these variables at four different values of $F$, with the mean removed. Before applying the DM algorithm, we normalize all variables to have unit variance. A bandwidth of $\epsilon=17$ for the Kernel (11) was found to be optimal to resolve the data manifold at all parameter values $F$ without being influenced by outliers.
Table 1: Spearman correlation of the top five Veros variables with the first two diffusion coordinates $\xi_{1}$ and $\xi_{2}$, where the $\xi_{n}$ were estimated at the four different values of the control parameter $F$. A shorthand notation encodes the meaning of the variables. E.g., ’rho deep TP’ is deep ocean density in the tropical Pacific. The first part (’rho’, ’salt’ or ’temp’) refers to density, salinity or temperature, respectively. The second part (’subs’ or ’deep’) refers to the subsurface ocean averaged until 1000m depth or the deep ocean below 1000m. The last part is geographical location, which can be the northern, tropical or southern Atlantic or Pacific (NA, TA, SA, NP, TP, SP), or the Southern Ocean (SO) and the Indian Ocean (IO). In bold text are key variables related to the mode of cold and fresh excursions in the deep north and tropical Atlantic directed towards the edge state. In italic are key variables related to the mode of fast cold excursions in the Southern ocean, which strongly increase the density in large parts of the deep ocean.
@@
To design a physical observables for EWS from $\xi_{1}$, the first option is again to use the Nyström extension based on all variables and interpolate the function $\xi_{1}$ to observations further back in time, in order to find evidence for increased fluctuations in the critical mode. One may also fit $\xi_{1}$ as a linear or non-linear function of the variables. In high dimensions it is sensible to only consider a subset of variables to find a parsimonious observable with the best signal-to-noise ratio when applied for EWS. We leave a treatment of this statistical optimization problem for future research, and consider here three simple examples. First, simply take the variable with highest correlation to $\xi_{1}$ as observable. This is north Atlantic deep ocean salinity (Fig. 11a-d), which indeed has the highest increase in variance of all individual features [15], increasing by a factor of 12.23 when going from $F=0.2957$ to $F=0.3557$. Second, fit $\xi_{1}$ to a linear combination of the two best features in Tab. 1 (at $F=0.3557$). This yields a variance increase by a factor of 14.21. Third, fit $\xi_{1}$ to a cubic polynomial of two variables, which are chosen as the two features among the best four (Tab. 1) that represent two different ocean sectors, thus giving some degree of spatial independence. This observable captures the directionality of the edge state well (Fig. 12), and shows an increase in variance by a factor of 12.60.
@@
Finally, $\xi_{1}$ can be mapped back into the physical space of full dimension, whereafter spatio-temporal anomalies from different time periods can be projected onto the mode. In particular, we propose to obtain the physical representation of this mode by averaging over the time points where $\xi_{1}$ is extremized. For instance, one can choose the data points with the top 5% largest and 5% smallest values of $\xi_{1}$. Averaging independently over these two sets of data points defines a positive and a negative phase of the mode. By taking the difference of the positive and negative patterns we obtain a pattern that describes the mode as a whole and that we can project data onto. This may be viewed as linear approximation that interpolates the physical mode linearly as a function of the value of $\xi_{1}$. Fig. 13 shows the first subdominant mode extracted in this way, comparing results far from (Fig. 13a,d) and close to the TP (Fig. 13b,e). The modes are projected down to the two-dimensional physical space of vertically averaged deep ocean of temperature and salinity. Far from the TP, the mode is characterized by a global cooling of the deep ocean initiated by abrupt (decadal-scale) cooling events in the Southern Ocean (seen as spikes of density increase in Fig. 11e-h), which are excited by noise in the multistability regime, as discussed in [15]. The resulting dense deep ocean water spreads throughout the deep ocean, before the signal decays on a multi-centennial time scale. There is no significant salinity signal.
@@
While these are encouraging results, in practice there is a risk of false positives and negatives, since here the critical physical mode only appears in its correct place at $\xi_{1}$ when already quite close to the TP. One needs to verify that the leading mode captured by $\xi_{1}$ is likely a critical mode. The guiding principle should be to look how strongly fluctuations in the leading diffusion coordinate, estimated at the current time slice, have increased compared to data slices back in time, and then set a level of statistical significance based on a reference period. In the Veros data one can in this way rule out the initially leading mode (Fig. 13a,d), because its variability does not keep increasing towards the TP (Fig. 11f-h), and the associated excursions do not last longer. But in principle there remains a chance for false positives if a new time scale separation would arise upon change of a control parameter for reasons not related to a TP.